consider two numbers. one of the numbers is 14 more than the other. the sum of their squares is 116. find the numbers
Answer by Alan3354(69443)

consider two numbers. one of the numbers is 14 more than the other. the sum of their squares is 116. find the numbers
----------
y = x + 14
y^2 + x^2 = 116
----
Sub for y in the 2nd equation.
(x+14)^2 + x^2 = 116
2x^2 + 28x + 196 = 116
x^2 + 14x + 40 = 0
(x+4)*(x+10) = 0
x = -4, y = 10
x = -10, y = 4